OpenStudy (anonymous):

y=x^2+1 find the vertex

OpenStudy (anonymous):

\[\Huge \frac{-b}{2a}\]

jimthompson5910 (jim_thompson5910):

y=x^2+1 really looks like y=1(x-0)^2+1 and it is in the form y = a(x-h)^2+k where (h,k) is the vertex We can see that h = 0 and k = 1 So the vertex is (0, 1)

OpenStudy (mathteacher1729):

y = x^2 + 1 is just y = x^2 but shifted vertically by +1. So the vertex of y = x^2 is just (0,0) and so the vertex of y = x^2 + 1 is (0,1). :)